\newcommand{\etalchar}[1]{$^{#1}$}
\begin{thebibliography}{GKLW18}

\bibitem[ABL14]{awasthi2014power}
Pranjal Awasthi, Maria~Florina Balcan, and Philip~M Long.
\newblock The power of localization for efficiently learning linear separators
  with noise.
\newblock In {\em Proceedings of the forty-sixth annual ACM symposium on Theory
  of computing}, pages 449--458. ACM, 2014.

\bibitem[AS18]{journals/corr/abs-1812-06369}
Emmanuel Abbe and Colin Sandon.
\newblock Provable limitations of deep learning.
\newblock {\em CoRR}, abs/1812.06369, 2018.

\bibitem[BDL18]{journals/corr/abs-1809-10787}
Digvijay Boob, Santanu~S. Dey, and Guanghui Lan.
\newblock Complexity of training relu neural network.
\newblock {\em CoRR}, abs/1809.10787, 2018.

\bibitem[BG15]{buet2015multinomial}
Francois Buet-Golfouse.
\newblock A multinomial theorem for hermite polynomials and financial
  applications.
\newblock {\em Applied Mathematics}, 6(06):1017, 2015.

\bibitem[BG17]{brutzkus2017globally}
Alon Brutzkus and Amir Globerson.
\newblock Globally optimal gradient descent for a convnet with gaussian inputs.
\newblock {\em arXiv preprint arXiv:1702.07966}, 2017.

\bibitem[BGS14]{conf/nips/BreslerGS14a}
Guy Bresler, David Gamarnik, and Devavrat Shah.
\newblock Structure learning of antiferromagnetic ising models.
\newblock In Zoubin Ghahramani, Max Welling, Corinna Cortes, Neil~D. Lawrence,
  and Kilian~Q. Weinberger, editors, {\em Advances in Neural Information
  Processing Systems 27: Annual Conference on Neural Information Processing
  Systems 2014, December 8-13 2014, Montreal, Quebec, Canada}, pages
  2852--2860, 2014.

\bibitem[DKS18]{DKS18a}
Ilias Diakonikolas, Daniel~M. Kane, and Alistair Stewart.
\newblock Learning geometric concepts with nasty noise.
\newblock In Ilias Diakonikolas, David~Kempe 0001, and Monika Henzinger,
  editors, {\em Proceedings of the 50th Annual ACM SIGACT Symposium on Theory
  of Computing, STOC 2018, Los Angeles, CA, USA, June 25-29, 2018}, pages
  1061--1073. ACM, 2018.

\bibitem[Fel16]{reference/algo/Feldman16b}
Vitaly Feldman.
\newblock Statistical query learning.
\newblock In {\em Encyclopedia of Algorithms}, pages 2090--2095. 2016.

\bibitem[FGV15]{journals/corr/FeldmanGV15}
Vitaly Feldman, Cristobal Guzman, and Santosh Vempala.
\newblock Statistical query algorithms for stochastic convex optimization.
\newblock {\em CoRR}, abs/1512.09170, 2015.

\bibitem[GKKT17]{goel2016reliably}
Surbhi Goel, Varun Kanade, Adam~R. Klivans, and Justin Thaler.
\newblock Reliably learning the relu in polynomial time.
\newblock In Satyen Kale and Ohad Shamir, editors, {\em Proceedings of the 30th
  Conference on Learning Theory, COLT 2017, Amsterdam, The Netherlands, 7-10
  July 2017}, volume~65 of {\em Proceedings of Machine Learning Research},
  pages 1004--1042. PMLR, 2017.

\bibitem[GKLW18]{journals/corr/abs-1810-06793}
Rong Ge, Rohith Kuditipudi, Zhize Li, and Xiang Wang.
\newblock Learning two-layer neural networks with symmetric inputs.
\newblock {\em CoRR}, abs/1810.06793, 2018.

\bibitem[GKM18]{conf/icml/GoelKM18}
Surbhi Goel, Adam~R. Klivans, and Raghu Meka.
\newblock Learning one convolutional layer with overlapping patches.
\newblock In Jennifer~G. Dy and Andreas~Krause 0001, editors, {\em ICML},
  volume~80 of {\em JMLR Workshop and Conference Proceedings}, pages
  1778--1786. JMLR.org, 2018.

\bibitem[GLM17]{journals/corr/abs-1711-00501}
Rong Ge, Jason~D. Lee, and Tengyu Ma.
\newblock Learning one-hidden-layer neural networks with landscape design.
\newblock {\em CoRR}, abs/1711.00501, 2017.

\bibitem[Kea98]{journals/jacm/Kearns98}
Michael~J. Kearns.
\newblock Efficient noise-tolerant learning from statistical queries.
\newblock {\em J. ACM}, 45(6):983--1006, 1998.

\bibitem[KK14]{klivans2014embedding}
Adam Klivans and Pravesh Kothari.
\newblock Embedding hard learning problems into gaussian space.
\newblock In {\em Approximation, Randomization, and Combinatorial Optimization.
  Algorithms and Techniques (APPROX/RANDOM 2014)}. Schloss
  Dagstuhl-Leibniz-Zentrum fuer Informatik, 2014.

\bibitem[KKKS11]{KKKS11}
Sham~M. Kakade, Adam Kalai, Varun Kanade, and Ohad Shamir.
\newblock Efficient learning of generalized linear and single index models with
  isotonic regression.
\newblock In {\em NIPS}, pages 927--935, 2011.

\bibitem[KKMS08]{kalai2008agnostically}
Adam~Tauman Kalai, Adam~R Klivans, Yishay Mansour, and Rocco~A Servedio.
\newblock Agnostically learning halfspaces.
\newblock {\em SIAM Journal on Computing}, 37(6):1777--1805, 2008.

\bibitem[KS09]{KalaiS09}
Adam Kalai and Ravi Sastry.
\newblock The isotron algorithm: High-dimensional isotonic regression.
\newblock In {\em COLT}, 2009.

\bibitem[MR18a]{journals/corr/abs-1810-04207}
Pasin Manurangsi and Daniel Reichman.
\newblock The computational complexity of training relu(s).
\newblock {\em CoRR}, abs/1810.04207, 2018.

\bibitem[MR18b]{MR2018}
Pasin Manurangsi and Daniel Reichman.
\newblock The computational complexity of training relu(s).
\newblock {\em CoRR}, abs/1810.04207, 2018.

\bibitem[Sha18]{Shamir18}
Ohad Shamir.
\newblock Distribution-specific hardness of learning neural networks.
\newblock {\em Journal of Machine Learning Research}, 19:32:1--32:29, 2018.

\bibitem[Sol17]{soltanolkotabi2017learning}
Mahdi Soltanolkotabi.
\newblock Learning relus via gradient descent.
\newblock In {\em Advances in Neural Information Processing Systems}, pages
  2007--2017, 2017.

\bibitem[SSSS17]{shalev2017failures}
Shai Shalev-Shwartz, Ohad Shamir, and Shaked Shammah.
\newblock Failures of gradient-based deep learning.
\newblock In {\em International Conference on Machine Learning}, pages
  3067--3075, 2017.

\bibitem[Sz{\"o}09]{szorenyi2009characterizing}
Bal{\'a}zs Sz{\"o}r{\'e}nyi.
\newblock Characterizing statistical query learning: simplified notions and
  proofs.
\newblock In {\em International Conference on Algorithmic Learning Theory},
  pages 186--200. Springer, 2009.

\bibitem[Val15]{valiant2015finding}
Gregory Valiant.
\newblock Finding correlations in subquadratic time, with applications to
  learning parities and the closest pair problem.
\newblock {\em Journal of the ACM (JACM)}, 62(2):13, 2015.

\bibitem[Ver]{vershynin2017four}
Roman Vershynin.
\newblock Four lectures on probabilistic methods for data science.

\bibitem[VW18]{vempala2018gradient}
Santosh Vempala and John Wilmes.
\newblock Gradient descent for one-hidden-layer neural networks: Polynomial
  convergence and sq lower bounds.
\newblock {\em arXiv preprint arXiv:1805.02677}, 2018.

\bibitem[Zha18]{Zhang18}
Chicheng Zhang.
\newblock Efficient active learning of sparse halfspaces.
\newblock In S{\'e}bastien Bubeck, Vianney Perchet, and Philippe Rigollet,
  editors, {\em Conference On Learning Theory, COLT 2018, Stockholm, Sweden,
  6-9 July 2018}, volume~75 of {\em Proceedings of Machine Learning Research},
  pages 1856--1880. PMLR, 2018.

\bibitem[ZSJ{\etalchar{+}}17]{zhong2017recovery}
Kai Zhong, Zhao Song, Prateek Jain, Peter~L Bartlett, and Inderjit~S Dhillon.
\newblock Recovery guarantees for one-hidden-layer neural networks.
\newblock {\em arXiv preprint arXiv:1706.03175}, 2017.

\bibitem[ZYWG18]{journals/corr/abs-1806-07808}
Xiao Zhang, Yaodong Yu, Lingxiao Wang, and Quanquan Gu.
\newblock Learning one-hidden-layer relu networks via gradient descent.
\newblock {\em CoRR}, abs/1806.07808, 2018.

\end{thebibliography}

\documentclass[letterpaper, 11pt]{article}

\usepackage[english]{babel}
\usepackage[utf8x]{inputenc}
\usepackage[T1]{fontenc}
\usepackage{chngcntr}
\usepackage{apptools}
\AtAppendix{\counterwithin{lemma}{section}}

\usepackage{geometry}
\usepackage{graphicx}
\usepackage{subfigure}
\usepackage{booktabs} 
\usepackage{amsfonts}       
\usepackage{authblk}

\usepackage{xcolor}
\usepackage{amsmath}
\usepackage{empheq}
\usepackage{amssymb}
\usepackage{amsthm}
\usepackage{comment}
\usepackage{bbm}
\usepackage{algorithm}
\usepackage[noend]{algpseudocode}
\newcommand{\E}{\mathbb{E}}
\newtheorem{theorem}{Theorem}
\newtheorem{lemma}{Lemma}
\newtheorem{claim}{Claim}
\newtheorem{fact}{Fact}
\newtheorem{corollary}{Corollary}

\newtheorem{definition}{Definition}

\newtheorem{assumption}{Assumption}
\usepackage{graphicx}
\newcommand{\D}{\mathcal{D}}
\newcommand{\calC}{\mathcal{C}}

\newcommand{\vx}{\textbf{x}}
\newcommand{\vw}{\textbf{w}}
\newcommand{\ve}{\textbf{e}}
\newcommand{\vz}{\textbf{z}}
\newcommand{\vv}{\textbf{v}}

\newcommand{\rmI}{\textbf{I}}
\newcommand{\relu}{\mathsf{ReLU}}
\newcommand{\sgn}{\mathsf{sign}}
\newcommand{\err}{\mathsf{err}}
\renewcommand{\hat}{\widehat}
\newcommand{\opt}{\mathsf{opt}}

\DeclareMathOperator*{\argmin}{arg\,min}
\allowdisplaybreaks

\title{Time/Accuracy Tradeoffs for Learning a ReLU with respect to
  Gaussian Marginals}

\author[]{Surbhi Goel\footnote{surbhi@cs.utexas.edu}}
\author[]{Sushrut Karmalkar\footnote{sushrutk@cs.utexas.edu}}
\author[]{Adam Klivans\footnote{klivans@cs.utexas.edu}}
\affil[]{Department of Computer Science, University of Texas at Austin}
\date{}
\begin{document}

\maketitle
\begin{abstract}
  We consider the problem of computing the best-fitting ReLU with
  respect to square-loss on a training set when the examples have been
  drawn according to a spherical Gaussian distribution (the labels can
  be arbitrary).  Let $\opt < 1$ be the population loss of the
  best-fitting ReLU.  We prove:

\begin{itemize}

\item Finding a ReLU with square-loss $\opt + \epsilon$ is as
  hard as the problem of learning sparse parities with noise, widely thought
  to be computationally intractable.  This is the first hardness
  result for learning a ReLU with respect to Gaussian marginals, and
  our results imply --unconditionally-- that gradient descent cannot
  converge to the global minimum in polynomial time.

\item There exists an efficient approximation algorithm for finding the
  best-fitting ReLU that achieves error $O(\opt^{2/3})$.  The
  algorithm uses a novel reduction to noisy halfspace learning with
  respect to $0/1$ loss. 

\end{itemize}
Prior work due to Soltanolkotabi \cite{soltanolkotabi2017learning} showed that gradient descent {\em can} find the best-fitting ReLU with respect to Gaussian marginals, if the training set is {\em exactly} labeled by a ReLU.

\end{abstract}

\section{Introduction}
A Rectified Linear Unit (ReLU) is a function parameterized by a weight vector $\vw \in \mathbb{R}^d$ that maps $\mathbb{R}^d \rightarrow \mathbb{R}$ as follows: $\relu_{\vw}(\vx) = \max(0, \vw \cdot \vx)$.  ReLUs are now the nonlinearity of choice in modern deep networks.  The computational complexity of learning simple neural networks that use the ReLU activation is an intensely studied area, and many positive results rely on assuming that the marginal distribution on the examples is a spherical Gaussian \cite{journals/corr/abs-1806-07808, journals/corr/abs-1711-00501, zhong2017recovery, journals/corr/abs-1810-06793}.  Recent work due to Soltanolkotabi \cite{soltanolkotabi2017learning} shows that gradient descent will learn a single ReLU in polynomial time, if the marginal distribution is Gaussian (see also \cite{brutzkus2017globally}).  His result, however, requires that the training set is {\em noiseless}; i.e., there is a ReLU that correctly classifies all elements of the training set.  

Here we consider the more realistic scenario of empirical risk minimization or learning a ReLU with noise (often referred to as {\em agnostically} learning a ReLU).  We assume that a learner has access to a training set from a joint distribution ${\cal D}$ on $\mathbb{R}^d \times \mathbb{R}$ where the marginal distribution on $\mathbb{R}^d$ is Gaussian but the distribution on the labels can be arbitrary within $[0,1]$.  We define $\opt = \min_{w, \|w\| \leq 1} \E_{x,y \sim {\cal D}}[(\relu_{w}(x) - y)^2]$, and the goal is to output a function of the form $\max(0, \vw \cdot \vx)$ with square-loss at most $\opt + \epsilon$. 

\subsection{Our Results}
Our main results give a trade-off between the accuracy of the output hypothesis and the running time of the algorithm.  We give the first evidence that there is no polynomial-time algorithm for finding a ReLU with error $\opt + \epsilon$, even when the marginal distribution is Gaussian:

\begin{theorem}[Informal version of Theorem \ref{thm:main-thm-lower-bound}] \label{thm:mainlower} Assuming hardness of the problem of learning sparse parities with noise, any algorithm for finding a ReLU on data drawn from a distribution with Gaussian marginals that has error at most $\opt + \epsilon$ runs in time $d^{\Omega(\log(1/\epsilon))}$.  \end{theorem}

Since gradient descent is known to be a {\em statistical-query} algorithm (see Section \ref{sec:sq}), a consequence of Theorem \ref{thm:mainlower} is the following:

\begin{corollary} Gradient descent fails to converge to the global minimum for learning the best-fitting ReLU with respect to square-loss in polynomial time, even when the marginals are Gaussian.  
\end{corollary}

This above corollary is unconditional (i.e. does not rely on any hardness assumptions) and shows the necessity of the realizable/noiseless setting in the work of Soltanolkotabi \cite{soltanolkotabi2017learning} and Brutzkus and Globerson \cite{brutzkus2017globally}. We also give the first approximation algorithm for finding the best-fitting ReLU with respect to Gaussian marginals:

\begin{theorem}[Informal version of Theorem \ref{thm:main-thm-upper-bound}]
There exists a polynomial-time algorithm for finding a ReLU with error $O(\opt^{2/3}) + \epsilon$.
\end{theorem}

The above result uses a novel reduction from learning a ReLU to the problem of learning a halfspace with respect to $0/1$ loss.   We note that the problem of finding a ReLU with error $O(\opt) + \epsilon$ remains an outstanding open problem. 

\subsection{Our Techniques}
\paragraph{Hardness Result.} For our hardness result, we follow the same approach as Klivans and Kothari \cite{klivans2014embedding} who gave a reduction from learning sparse parity with noise to the problem of agnostically learning {\em halfspaces} with respect to Gaussian distributions.  The idea is to embed examples drawn from $\{-1,1\}^{d}$ into $\mathbb{R}^{d}$ by multiplying each coordinate with a random draw from a half-normal distribution.  The key technical component in their result is a correlation lemma showing that for a parity function on variables indicated by index set $S$, the majority function on the same index set is weakly correlated with a {\em Gaussian lift} of the parity function on $S$.

In our work we must overcome two technical difficulties.  First, in the Klivans and Kothari result, it is obvious that for distributions induced by learning sparse parity with noise, the best fitting majority function will be the one that is defined on inputs specified by $S$.  In our setting with respect to ReLUs, however, the constant function $1/2$ will have square-loss $1/4$, and this may be much lower than the square-loss of any function of the form $\max(0, \vw \cdot \vx)$.  Thus, we need to prove the existence of a gap between the correlation of ReLUs with random noise (see Claim \ref{claim:inactive_index}) versus the correlation of ReLUs with parity (see Claim \ref{claim:active_index}).    

Second, Klivans and Kothari use known formulas on the discrete Fourier coefficients of the majority function and an application of the central limit theorem to analyze how much the best-fitting majority correlates with the Gaussian lift of parity.  No such bounds are known, however, for the ReLU function.  As such we must perform a (somewhat involved) analysis of the ReLU function's Hermite expansion in order to obtain quantitative correlation bounds.

\paragraph{Approximation Algorithm.} For our polynomial-time algorithm that outputs a ReLU with error $\mathsf{O(\opt^{2/3})} + \epsilon$, we apply a novel reduction to agnostically learning halfspaces.  We give a simple transformation on the training set to a Boolean learning problem and show that the weight vector $\vw$ corresponding to the best fitting {\em halfspace} on this transformed data set is not too far from the weight vector corresponding to the best fitting ReLU.  We can then apply recent work for agnostically learning halfspaces with respect to Gaussians that have constant-factor approximation error guarantees.  The exponent $2/3$ appears due to the use of an averaging argument (see Section \ref{sec:approx}).

\subsection{Related Work}
Several recent works have proved hardness results for finding the best-fitting ReLU with respect to square loss (equivalently, agnostically learning a ReLU with respect to square loss).  Results showing NP-hardness (e.g.,  \cite{journals/corr/abs-1810-04207,journals/corr/abs-1809-10787}) use marginal distributions that encode hard combinatorial problems.  The resulting marginals are far from Gaussian.  Work due to Goel et al. \cite{goel2016reliably} uses a reduction from sparse parity with noise but only obtains hardness results for learning with respect to discrete distributions (uniform on $\{0,1\}^{d}$).  

Using parity functions as a source of hardness for learning deep networks has been explored recently by Shalev-Shwartz et. al.~\cite{shalev2017failures} and Abbe and Sandon~\cite{journals/corr/abs-1812-06369}.  Their results, however, do not address the complexity of learning a {\em single} ReLU or consider the case of Gaussian marginals.  Shamir~\cite{Shamir18} proved that gradient descent fails to learn certain classes of neural networks with respect to Gaussian marginals, but these results do not apply to learning a single ReLU~\cite{vempala2018gradient}.

In terms of positive results for learning a ReLU, work due to Kalai and Sastry \cite{KalaiS09} (and follow-up work \cite{KKKS11}) gave the first efficient algorithm for learning any generalized linear model (GLM) that is monotone and Lipschitz, a class that includes ReLUs.  Their algorithms work for {\em any} distribution and can tolerate bounded, mean-zero and additive noise. Soltanolkotabi \cite{soltanolkotabi2017learning} and Brutzkus and Globerson \cite{brutzkus2017globally} were the first to prove that gradient descent converges to the unknown ReLU in polynomial time with respect to Gaussian marginals as long as the labels have {\em no noise}. Other works for learning one-layer ReLU networks with respect to Gaussian marginals or marginals with milder distribution assumptions \cite{journals/corr/abs-1806-07808, journals/corr/abs-1711-00501, zhong2017recovery, journals/corr/abs-1810-06793, conf/icml/GoelKM18, MR2018} also assume a noiseless training set or training set with mean-zero i.i.d. (typically sub-Gaussian) noise. This is in contrast to the setting here (agnostic learning), where we assume nothing about the noise model. 

There are several works for the related (but different) problem of agnostically learning {\em halfspaces} with respect to Gaussian marginals \cite{kalai2008agnostically, awasthi2014power, Zhang18, DKS18a}.  While agnostically learning ReLUs may seem like an easier problem than agnostically learning halfspaces (at first glance the learner sees ``more information'' from the ReLU's real-valued labels),  the quantitative relationship between the two problems is still open.  In the halfspace setting, we can assume without loss of generality that an adversary has flipped an $\opt$ fraction of the labels.  In contrast, in the setting with ReLUs and square loss, it is possible for the adversary to corrupt {\em every} label.

\section{Preliminaries}
Define $\relu(a) = \max(0,a)$ and the set of functions $\calC_\relu:= \{\relu_\vw| \vw \in \mathbb{R}^d, \| \vw \|_2 \leq 1\}$ where $\relu_\vw(\vx) = \max(0, \vw \cdot \vx)$. Define $\sgn(a)$ to be 1 if $a \ge 0$ and -1 otherwise. Let $\err_{\D}(h):= \E_{(\vx,y) \sim \D}[(h(\vx) - y)^2]$, Also define $\opt_\D(\calC) = \min_{c \in \calC} \err_{\D}(c)$ to be the error of the best-fitting $c \in \calC$ for distribution $\D$. We will use $\vx_{-i}$ to denote the vector $\vx$ restricted to the indices except $i$. The `half-normal' distribution will refer to the standard normal distribution truncated to $\mathbb{R}^{\geq 0}$. We will use $n$ and its subscripted versions to denote natural numbers unless otherwise stated. In this paper, we will suppress the confidence parameter $\delta$, since one can use standard techniques to amplify the probability of success of our learning algorithms.

\paragraph{Agnostic learning.} The model of learning we work with in the paper is the agnostic model of learning. In this model the labels are allowed to be arbitrary and the task of the learner is to output a hypothesis within an $\epsilon$ error of the optimal. More formally,
\begin{definition}
A class $\calC$ is said to be agnostically learnable in time $t$ over the Gaussian distribution to error $\epsilon$ if there exists an algorithm ${\cal A}$ such that for any distribution $\D$ on $X \times Y$ with the marginal on $X$ being Gaussian, ${\cal A}$ uses at most $t$ draws from ${\cal D}$, runs in time at most $t$, and outputs a hypothesis $h \in \calC$ such that $\err_\D(h) \le  \opt_\D(\calC) + \epsilon$.
\end{definition}
We assume that ${\cal A}$ succeeds with constant probability.  Note that the algorithm above outputs the ``best-fitting'' $c\in \calC$ with respect to $\D$ up to an additive $\epsilon$.  We will denote $\widehat{\err}_{\mathcal{S}}(h)$ to be the empirical error of $h$ over samples $\mathcal{S}$.

\paragraph{Learning Sparse Parities with Noise.} In this work we will show that agnostically learning $\calC_\relu$ over the Gaussian distribution is as hard as the problem of learning sparse parities with noise over the uniform distribution on the hypercube.
\begin{definition}[$k$-SLPN]
Given access to samples drawn from the uniform distribution over $\{\pm 1\}^d$ and target function $y$ being the parity function over an unknown set $S \subseteq [d]$ of size $k$, the problem of learning sparse parities with noise is the problem of recovering the set $S$ given access to noisy labels where the label is flipped with probability $\eta$.
\end{definition}
Learning sparse parities with noise is generally considered to be a computationally hard problem and has been used to give hardness results for both supervised \cite{goel2016reliably} and unsupervised learning problems \cite{conf/nips/BreslerGS14a}.  The current best known algorithm for solving sparse parities with constant noise rate is due to Valiant~\cite{valiant2015finding} and runs in time $\approx d^{0.8k}$.

\begin{assumption}\label{conj:SLPN}
Any algorithm for solving $k$-SLPN up to constant error must run in time $d^{\Omega(k)}$.
\end{assumption}

\paragraph{Gaussian Lift of a Function} Our reduction will require the following definition of a Gaussian lift of a boolean function from \cite{klivans2014embedding}. 

\begin{definition}[Gaussian lift \cite{klivans2014embedding}]
	The Gaussian lift of a function $f : \{ \pm 1 \}^d \rightarrow \mathbb{R}$ is the function $f^{\gamma} : \mathbb{R}^d \rightarrow \mathbb{R}$ such that for any $x \in \mathbb{R}^d$, $f^{\gamma}(x) = f(\sgn(x_1), \dots, \sgn(x_d))$. 
\end{definition} 

\paragraph{Hermite Analysis and Gaussian Density}
We will assume that the marginal over our samples $\vx$ is the standard normal distribution $N(0, I_{d})$. This implies that $\vw \cdot \vx$ for a vector $\vw$ is distributed as $N(0, \|\vw\|^2)$. We recall the basics of Hermite analysis. We say a function $f: \mathbb{R} \rightarrow \mathbb{R}$ is square integrable if $\E_{N(0, 1)}[f^2] < \infty$.  For any square integrable function $f$ define its Hermite expansion as $f(x) = \sum_{i=0}^\infty \hat{f}_i \bar{H}_i(x)$ where $\bar{H}_i(x) = \frac{H_i(x)}{\sqrt{i!}}$ are the normalized Hermite polynomials, and $H_i$ the unnormalized (probabilists) Hermite polynomials. The normalized Hermite polynomials form an orthonormal basis with respect to the univariate standard normal distribution ($\E[\bar{H}_i(x)\bar{H}_j(x)] =\delta_{ij}$).   The associated inner product for square integrable functions $f, g : \mathbb{R} \rightarrow \mathbb{R}$ is defined as $\langle f, g \rangle := \E_{x \sim N(0, 1)} [ f(x)g(x)]$. Each coefficient $\hat{f}_i$ in the expansion of $f(x)$ satisfies $\hat{f}_i = E_{x \sim N(0,1)}[f(x) \bar{H}_i(x)]$.

  We will need the following facts about Hermite polynomials. 
\begin{fact}
	For all $m \ge 0$, $H_{2m+1}(0) = 0$ and $H_{2m} = (-1)^m\frac{(2m)!}{m!2^m}$.
\end{fact}
\begin{fact}[\cite{kalai2008agnostically}]\label{lem:hsgn}
	$\hat{\sgn}_0 = 0$ and for $i \ge 1$, $\hat{\sgn}_i = \sqrt{\frac{2}{\pi i!}}H_{i-1}(0)$.
\end{fact}

\section{Hardness of Learning ReLU}
In this section, we will show that if there is an algorithm that agnostically learns a ReLU in polynomial time, then there is an algorithm for learning sparse parities with noise in time $d^{o(k)}$, violating Assumption \ref{conj:SLPN}. We will follow the approach of \cite{klivans2014embedding}. Let $\chi_S$ be an unknown parity for some $S \subseteq [d]$. We will show that there is an unbiased ReLU that is correlated with the Gaussian lift of the unknown sparse parity function.  Notice that dropping a coordinate $j \in S$ from the input samples makes the labels of the resulting training set totally independent from the input.  In contrast, dropping $j \notin S$ results in a training set that is still labeled by a noisy parity. Therefore, we can use an agnostic learner for ReLUs to detect a correlated ReLU and distinguish between the two cases.  This allows us to identify the variables in $S$ one by one. 

We formalize the above approach by first proving the following key property,
\begin{lemma}[$\relu$ Correlation Lemma]\label{thm:relu_correlated}
	Let $\chi_{S}^{\gamma}$ denote the Gaussian lift of the parity on variables in $S \subset [d]$. For every $S \subset [d]$ with $|S| \leq k$ and $k = 4l + 2$ for some $l \ge 0$, there exists $\relu_{\vw_S}$ such that $\langle \relu_{\vw_S}, \chi_{\alpha}^{\gamma} \rangle \ge 2^{-O(k)}$ where $\relu_{\vw_S}$ only depends on variables in $S$. 
\end{lemma}
\begin{proof} 
	Let $\vw_S = \frac{1}{\sqrt{2\pi k}} \sum_{i\in S} \ve^{(i)}$ where $\ve^{(i)}$ is 1 at coordinate $i$ and 0 everywhere else. We will show that
    \begin{align*}
    \langle \relu_{\vw_S}, \chi^{\gamma}_S \rangle = \frac{1}{\sqrt{2\pi}}\E_{\vz \sim \mathcal{N}(0, \rmI_d)}\left[\relu\left(\frac{\sum_{i \in S} z_i}{\sqrt{k}}\right)\left(\prod_{i \in S}\sgn(z_i)\right)\right] \ge 2^{-O(k)}.
    \end{align*}

	Let $\widehat{\sgn}_{n}$ and $\widehat{\relu}_n$ denote the degree $n$ Hermite coefficients of the $\sgn$ function and $\relu$ function respectively. It is easy to see that the Hermite expansion of the Gaussian lift of a parity supported on $S$ is, 
	\begin{equation}\label{eqn:h_expansion_parity}
	\chi_{S}^{\gamma}(\vz) = \prod_{i \in S}\sgn(z_i) = \prod_{i \in S}\left(\sum_{n=0}^{\infty} \hat{\sgn}_n \bar{H}_n(z_i) \right) = \sum_{n_1, \ldots, n_k}\prod_{i \in S} \hat{\sgn}_{n_i} \bar{H}_{n_i}(z_i)	\end{equation}

In order to finish the proof of Lemma~\ref{thm:relu_correlated} we will need the expansion of $\relu\left(\frac{\sum_i z_i}{\sqrt{k}}\right)$ in terms of products of univariate Hermite polynomials. Toward this end we establish the following claims.
\begin{claim}[Hermite expansion: univariate $\relu$] \footnote{Proof in supplementary material}\label{lem:hrelu}
	$\hat{\relu}_0 = 1/\sqrt{2\pi}$, $\hat{\relu}_1 = 1/2$ and for $i \ge 2$, $\hat{\relu}_i = \frac{1}{\sqrt{2\pi i!}}(H_{i}(0) + iH_{i-2}(0))$.
\end{claim}
	\begin{claim}[Hermite expansion: multivariate $\relu$] \label{claim:h_expansion_relu} For any $S \subseteq [d]$ with $|S| = k$,
	\[ 	\relu\left(\frac{\sum_{i \in S} z_i}{\sqrt{k}}\right) = \sum_{n=0}^{\infty} \frac{\hat{\relu}_n}{k^{n/2}}\cdot\sum_{n_1 + \ldots + n_k = n}\left(\frac{n!}{n_1! \cdots n_k!}\right)^{1/2}\prod_{j=1}^k\bar{H}_{n_j}(z_j) \]
	\end{claim}
Combining Equation~\ref{eqn:h_expansion_parity} and Claim~\ref{claim:h_expansion_relu} now yields,
\begin{align*}
&\E_{\vz \sim \mathcal{N}(0, \rmI_d)}\left[\relu\left(\frac{\sum_{i \in S} z_i}{\sqrt{k}}\right)\cdot\prod_{i \in S}\sgn(z_i)\right] \\
&= \E_{\vz \sim \mathcal{N}(0, \rmI_d)}\left[\left(\sum_{n=0}^{\infty} \frac{\hat{\relu}_n}{k^{n/2}}\sum_{n_1 + \ldots + n_k = n}\left(\frac{n!}{n_1! \cdots n_k!}\right)^{1/2}\prod_{i=1}^k \bar{H}_{n_i}(z_i)\right)\right.\\
&\qquad \left. \times\left(\sum_{m_1, \ldots, m_k}\prod_{j=1}^k \hat{\sgn}_{m_j} \bar{H}_{m_j}(z_j)\right)\right]\\
&= \sum_{n=0}^{\infty} \frac{\hat{\relu}_n}{k^{n/2}}\sum_{n_1 + \ldots + n_k = n}\sum_{m_1, \ldots, m_k}\left(\frac{n!}{n_1! \cdots n_k!}\right)^{1/2}\prod_{i=1}^k\hat{\sgn}_{m_i}\E[\bar{H}_{n_i}(z_i) \bar{H}_{m_i}(z_i)]\\
&= \sum_{n=0}^{\infty} \frac{\hat{\relu}_n}{k^{n/2}}\sum_{n_1 + \ldots + n_k = n}\left(\frac{n!}{n_1! \cdots n_k!}\right)^{1/2}\prod_{i=1}^k \hat{\sgn}_{n_i}\\
\end{align*}
 From Fact \ref{lem:hsgn} and Claim \ref{lem:hrelu} we see that $\hat{\sgn}_{2m} = 0$ and $\hat{\relu}_{2m + 1} = 0$ for $m \ge 1$. Additionally, since $\hat{\sgn}_0 = 0$ we see that each $n_i \geq 1$. This gives us, 
\begin{align*}
& \E_{\vz \sim \mathcal{N}(0, \rmI_d)}\left[\relu\left(\frac{\sum_{i \in S} z_i}{\sqrt{k}}\right)\cdot\prod_{i \in S}\sgn(z_i)\right] \\
&= \sum_{n = k}^{\infty} \frac{1}{\sqrt{2 \pi n!}k^{n/2}}(H_n(0) + nH_{n-2}(0))\sum_{\substack{n_1, \ldots, n_k \ge 1\\ n_1 + \ldots + n_k = n}}\left(\frac{n!}{n_1! \cdots n_k!}\right)^{1/2}\prod_{i=1}^k \sqrt{\frac{2}{\pi n_j!}}H_{n_j - 1}(0)\\
&= \sum_{n = k}^{\infty} \frac{1}{\sqrt{2\pi} k^{n/2}}(H_n(0) + nH_{n-2}(0))\sum_{\substack{n_1, \ldots, n_k \ge 1\\ n_1 + \ldots + n_k = n}}\left(\frac{1}{n_1! \cdots n_k!}\right)^{3/2}\prod_{i=1}^k \sqrt{\frac{2}{\pi}} H_{n_j - 1}(0).
\end{align*}
To finish the proof of Lemma~\ref{thm:relu_correlated}, we will look at each term in the outer summation above. Let the term for any fixed $n \ge k$ be denoted by $T_n$. Since $\bar{H}_{i}(0) = 0$ for odd $i$, observe that $T_n$ is non-zero if and only if $n$ is even and each $n_i = 2n^\prime_i + 1$ for $n^\prime_i \ge 0$. We have
\begin{align*}
T_n &= \frac{1}{\sqrt{2\pi}k^{\frac{n}{2}}}\left((-1)^{\frac{n}{2}}\frac{n!}{(n/2)!2^{\frac{n}{2}}} + n (-1)^{\frac{n}{2}-1}\frac{(n - 2)!}{(\frac{n}{2}-1)!2^{\frac{n}{2}-1}}\right) \\
&\qquad \times \sum_{\substack{n^\prime_i \ge 0 \\ \sum_j n^\prime_j = \frac{n-k}{2}}}\left(\frac{1}{(2n^\prime_1 + 1)! \cdots (2n^\prime_k +1)!}\right)^{3/2}\prod_{j=1}^k \sqrt{\frac{2}{\pi}}(-1)^{n^\prime_j}\frac{(2n^\prime_j)!}{n^\prime_j!2^{n^\prime_j}}\\
&= \frac{(-1)^{\frac{n}{2} -1} n!}{\sqrt{2\pi}k^{\frac{n}{2}}(n/2)!2^{\frac{n}{2}}}\left(-1 + \frac{n}{n - 1}\right)\frac{2^{\frac{k}{2}}(-1)^{\frac{n - k}{2}}}{\pi^{\frac{k}{2}}}\\
&\qquad \times \sum_{\substack{n^\prime_j \ge 0\\ \sum_j n^\prime_j = \frac{n-k}{2}}}\left(\frac{1}{(2n^\prime_1 + 1)! \cdots (2n^\prime_k +1)!}\right)^{3/2}\frac{(2n^\prime_1)!\cdots (2n^\prime_k)!}{n^\prime_1! \cdots n^\prime_k!}\\
&= \frac{(-1)^{1 + \frac{k}{2}} n!}{\sqrt{2\pi}k^{\frac{n}{2}}(n - 1)(n/2)!2^{\frac{n - k}{2}}\pi^{\frac{k}{2}}}\sum_{\substack{n^\prime_i \ge 0\\ \sum_j n^\prime_j = \frac{n - k}{2}}}\left(\frac{1}{(2n^\prime_1 + 1)! \cdots (2n^\prime_k +1)!}\right)^{3/2}\frac{(2n^\prime_1)!\cdots (2n^\prime_k)!}{n^\prime_1! \cdots n^\prime_k!}
\end{align*}
Since $k = 4l + 2$ (by assumption), $T_n > 0$ for all even $n \ge k$ and equal to 0 for all odd $n$. Thus $\sum_{n = k}^\infty T_n > T_k$. Lower bounding $T_k$, we have
\begin{align*}
T_{k} &= \frac{(4l + 2)!}{\sqrt{2\pi}(4l + 2)^{2l + 1}(4l + 1)(2l+1)!\pi^{2l + 1}}\\
&\approx \frac{1}{\sqrt{2\pi}(4l + 2)^{2l + 1}(4l + 1)\pi^{2l + 1}} \sqrt{2 \pi (4l+2)} \left(\frac{4l + 2}{e}\right)^{4l + 2} \frac{1}{\sqrt{2 \pi (2l + 1)}}\left(\frac{e}{2l + 1}\right)^{2l + 1}\\
&= \frac{1}{\sqrt{\pi}(4l + 1)} \left(\frac{2}{e\pi}\right)^{2l + 1} = 2^{-O(k)}
\end{align*}
\end{proof}

Now we present our main algorithm (Algorithm~\ref{algo:main}) that reduces learning sparse parities with noise to agnostically learning ReLUs and a proof of its correctness.
\begin{algorithm}
	\caption{Learning Sparse Parities with Noise using Agnostic ReLU learner}\label{algo:main}
	\hspace*{\algorithmicindent} \textbf{Input} Training set $\mathcal{S}$ of $M_1$ samples $(\vx^{i}, y^{i})_{i=1}^{M_1}$, validation set $\mathcal{V}$ of $M_2$ samples\\
     \hspace*{\algorithmicindent} \quad \quad \quad $(\vx^{i}, y^{i})_{i=M_1 + 1}^{M_1 + M_2}$, error parameter $\epsilon$, Agnostic ReLU learner $\mathcal{A}$  \\
	\hspace*{\algorithmicindent} \textbf{Output} Set of relevant variables $V_{rel}$
	\begin{algorithmic}[1]
		\State Set $V_{rel} = \emptyset$
		\State Set $\mathcal{S}_{1}, \ldots, \mathcal{S}_{d} := \emptyset$
		\For{$i = 1$ to $M_1 + M_2$}
		\State Draw $n$ independent univariate half Gaussians $g_1, \ldots, g_d$
		\State Construct $\vx^\prime$ such that for all $j \in [d]$, $x^\prime_j := g_j x^{i}_j$ and set $y^\prime = \frac{y^{i} + 1}{2}$
        \State For all $j \in [d]$, if $i \le M_1$ add $(\vx^\prime_{-j}, y^\prime)$ to $\mathcal{S}_j$ else to $\mathcal{V}_j$
		\EndFor 
		\For{$j \in [d]$}
		\State Run $\mathcal{A}$ on $\mathcal{S}_j$ to obtain hypothesis $h_j$
		\State Compute $\hat{\err}_{\mathcal{V}_j}(h_j)$
		\If {$\hat{\err}_{\mathcal{V}_j}(h_j) \ge \frac{1}{2} - \frac{1}{4 \pi} - \epsilon/4$}
		\State Add $j$ to $V_{rel}$
		\EndIf
		\EndFor
		\State Return $V_{rel}$
	\end{algorithmic}
\end{algorithm}
\begin{theorem} \label{thm:main-thm-lower-bound} 
If there is an algorithm to agnostically learn unbiased ReLUs on the Gaussian distribution in time and samples $T(d, 1/\epsilon)$, then there is an algorithm to solve $k$-SLPN in time $O\left(\frac{2^{O(k)}}{(1 - 2\eta)^2}\log(d)\right) + O(d)T\left(d, \frac{2^{O(k)}}{1 - 2\eta}\right)$ where $\eta$ is the noise rate.

In particular, if Assumption 1 is true, then any algorithm for agnostically learning (unbiased) ReLUs on the Gaussian distribution must run in time $d^{\Omega(\log(1/\epsilon))}$.
\end{theorem}
\begin{proof}
	Given a set of samples from the $k$-SPLN problem, we claim that Algorithm~\ref{algo:main} can recover all indices $j$ belonging to the sparse parity when run with appropriate parameters. 
    We will first show that if a variable is relevant then the error is smaller compared to when it is irrelevant. It is easy to see that $y^\prime$ is $\frac{\prod_{i \in S}\sgn(x^\prime_i) + 1}{2}$ with probability $1 - \eta$ and $\frac{1 - \prod_{i \in S}\sgn(x^\prime_i)}{2}$ otherwise. Let $\D_j$ denote the distribution obtained by dropping the $j$th coordinate from the lifted distribution and let $S$ denote the set of active indices of the parity. The proof of the theorem follows from the following claims,
		
	\begin{claim} \label{claim:inactive_index}
	If $j \in S$ then for all $\vw$, $\err_{\D_j}(\relu_\vw) = \frac{\|\vw\|^2}{2} - \frac{\|\vw\|}{\sqrt{2\pi}} + \frac{1}{2} \geq \frac{1}{2} - \frac{1}{4\pi}$.
	\end{claim}
	\begin{claim} \label{claim:active_index}
	If $j \notin S$ then there exists $\vw^*$ with $\|\vw^*\| = \frac{1}{\sqrt{2 \pi}}$ such that 
	$\err_{\D_j}(\relu_{\vw^*}) < \frac{1}{2} - \frac{1}{4\pi} - \frac{2^{-O(k)}}{1 - 2 \eta}.$
	\end{claim}	
    Claims~\ref{claim:inactive_index} and ~\ref{claim:active_index} imply that  we have a gap of at least $\frac{2^{-O(k)}}{1 - 2 \eta} = \frac{2^{-ck}}{1 - 2 \eta}$ for some $c> 0$ between the relevant and irrelevant variable case. Setting $\epsilon =\frac{2^{-ck}}{1 - 2 \eta}$ in Algorithm \ref{algo:main} will let us detect this gap.  Since $\mathcal{A}$ is an agnostic learner for ReLU, as long as $M_1 = T(d, 2/\epsilon)$ we know that with probability $2/3$, for all $j \in S$, $\mathcal{A}$ runs on $\mathcal{S}_j$ and outputs $h_j$ such that $\err_{\D_j}(h_j) \le \min_{\vw} \err_{\D_j}(\relu_\vw) \le \frac{1}{2} - \frac{1}{4\pi} - \epsilon/2$, and for all $j \notin S$, $\err_{\D_j}(h_j) \ge \frac{1}{2} - \frac{1}{4 \pi}$.

    Using standard concentration inequalities for sub-Gaussian and subexponential random variables \cite{vershynin2017four} we see that using a validation set of $M_2 = 100/\epsilon^2$ samples, we have for all $j$, $|\hat{\err}_{\mathcal{V}_j}(h_j) - \err_{\D_j}(h_j)| \le \epsilon/4$. Therefore, we can differentiate the two cases as in the Algorithm with confidence $> 1/2$. It is easy to see that the run time of the algorithm is $O(d) T(d, 2/\epsilon) +O(1/\epsilon^2)$, and that this can be amplified to obtain an algorithm with any desired confidence using standard techniques.
\end{proof}

\section{Lower Bounds for SQ Algorithms} \label{sec:sq}
A consequence of Theorem~\ref{thm:main-thm-lower-bound} is that any {\em statistical-query} algorithm for agnostically learning a ReLU with respect to Gaussian marginals yields a statistical-query algorithm for learning parity functions on $k$ unknown input bits. This implies that there is no polynomial time statistical-query (SQ) algorithm that learns a ReLU with respect to Gaussian marginals for a certain restricted class of queries. 

\paragraph{SQ Model.} A SQ algorithm is a learning algorithm that succeeds given only estimates of $\E_{\vx,y \sim {\cal D}}[q(\vx,y)]$ for query functions of the learner's choosing to an oracle up to a fixed tolerance parameter (see for example \cite{journals/jacm/Kearns98,reference/algo/Feldman16b}).  We restrict ourselves to queries that are either {\em correlation} queries, that is, $\E_{\vx,y \sim {\cal D}}[y \cdot g(\vx)]$ for any function $g$, or queries that are {\em independent of the target}, that is, $\E_{\vx,y \sim {\cal D}}[h(\vx)]$ for any function $h$. For example, the $i^{th}$ coordinate of the gradient with respect to $\vw$ of the quantity $(\relu_{\vw}(\vx) - y)^2$, i.e. $2 \cdot 1_{+} (\vw \cdot \vx) \cdot (\relu_{\vw}(\vx)  - y) \cdot \vx_i$ can be simulated by a correlation query $g(\vx) = -2 \cdot 1_{+} (\vw \cdot \vx) \cdot \vx_i$ and query $h(\vx) = 2 \cdot 1_{+} (\vw \cdot \vx) \cdot \relu_{\vw}(\vx) \cdot \vx_i$ independent of the target. Therefore queries with sufficiently small tolerance allow us to simulate gradient descent for the loss function $L(\vw) = \E_{\vx,y \sim {\cal D}}[(\relu_{\vw}(\vx) - y)^2]$ \footnote{Feldman et. al. \cite{journals/corr/FeldmanGV15} have shown that a broad class of first order convex optimization methods including gradient descent -- but excluding stochastic gradient descent -- can be simulated using statistical queries}. 

\paragraph{SQ Dimension.} We define the inner product of two functions $f(\vx), g(\vx)$ with respect to a distribution $\D$ to be $\langle f, g \rangle_{\D} := \E_{\vx \sim \D}[f(\vx) g(\vx)]$. The norm of a function $f$ is just $\sqrt{\langle f, f \rangle_{\D}}$ and two functions $f \ne g$ are said to be orthogonal if $\langle f, g \rangle_{\D} = 0$. The SQ dimension of a function class $\mathcal{F}$ is the largest number of pairwise orthogonal functions that belong to the function class. The following theorem from \cite{szorenyi2009characterizing} gives a lower bound on the number of statistical queries needed to learn the function class in terms of its SQ dimension.

\begin{theorem}[Restatement of Theorem from \cite{szorenyi2009characterizing}]\label{thm:sq-szorenyi-proof} 
	Let $\mathcal{F}$ be a concept class and let $s$ be the SQ dimension of $\mathcal{F}$ with respect to $\D$. Then any learning algorithm that uses tolerance parameter lower bounded by $\tau > 0$ and has access to an oracle that returns $\tau$-approximate expectations (with respect to $\D$) of unit norm correlation queries and queries that are independent of the target, requires at least $(s\tau^2 -1)/2$ queries. 
\end{theorem}

In this model of learning, we show the following lower bound for the problem of learning ReLUs over the Gaussian distribution.
\begin{theorem}\label{thm:sq-lower-bound} 
	Any SQ algorithm for agnostically learning a ReLU with respect to any distribution $\D$ satisfying Gaussian marginals over the attributes, requires $d^{\Omega(\log(1/\epsilon))}$ unit norm correlation queries or queries independent of the target with tolerance $\frac{1}{\text{poly}(d, 1/\epsilon)}$  to an oracle that returns $\tau$-approximate expectations with respect to $\D$.
\end{theorem}
\begin{proof} 
Define the problem of `restricted $k$-sparse parities' as the problem of learning an unknown parity function $\chi_S$ over set $S$, where $S$ contains $k$ out of the first $d$ variables over $\D$ with input distribution $\D_b \times \mathcal{N}(0, I_d)_+$. Here $\D_b$ is the uniform distribution on $\{\pm 1\}^d$ and the labels $y(\vx)$ are given by $\chi_S(\vx)$. It is easy to see that Theorem~\ref{thm:sq-szorenyi-proof} implies that we require $d^{\Omega(k)}$ unit norm queries to learn this function class from queries to an oracle $\mathcal{O}$ with tolerance $1/\text{poly}(d, 2^k)$. 

We give a proof by contradiction. Suppose we can agnostically learn ReLUs with respect to Gaussian marginals using an SQ algorithm ${\cal A}$ with $d^{o(\log \frac{1}{\epsilon})}$ queries to the corresponding oracle with tolerance $1/\text{poly}(d, \frac{1}{\epsilon})$. We will show how to use $\mathcal{A}$ to design an SQ algorithm for the problem of learning restricted $k$-sparse parities using $d^{o(k)}$ queries contradicting Theorem \ref{thm:sq-szorenyi-proof}.

We essentially use the same approach as in Theorem~\ref{thm:main-thm-lower-bound}. In order to learn the parity function, the reduction requires us to use an SQ learner to solve $d$ different ReLU regression problems. Consider the mapping $\nu$ that takes as input $\vx = (x_1, \dots x_d, g_1, \dots, g_d)$ and returns $\vx' = (x_1 g_1, \dots, x_d g_d)$. Using this transformation, the ReLU regression problems we need to minimize are $\E_{(\vx, y) \sim \D}\left[\left(\relu_{\vw}(\nu(\vx)_{-i}) - \frac{y + 1}{2}\right)^2\right]$ up to an additive $\epsilon = 2^{-ck}$. Observe that for $(\vx, y) \sim \D$, $(\vx', y') = (\nu(\vx), (y+1)/2)$ is distributed according to some $\D^{\prime}$ where the distribution on $\nu(\vx)$ is $\mathcal{N}(0, I_d)$. Thus we can use $\mathcal{A}$ on this distribution to solve the optimization problem. However, in order to run $\mathcal{A}$, we need to simulate the queries $\mathcal{A}$ asks its oracle using $\mathcal{O}$. To do so, for any correlation query function $g$ that ${\cal A}$ chooses, that is, query $\E_{(\vx', y') \sim \D^\prime}\left[g(\vx') \cdot y'\right]$ we use correlation query function $g' = \frac{1}{2} g \circ \nu$ to $\mathcal{O}$ and for any query $h$ that ${\cal A}$ chooses that is independent of the target, we  query function $h' = \frac{1}{2} h \circ \nu$.

Since for $k = \Theta(\log \frac{1}{\epsilon})$, such an algorithm would solve the problem of `restricted $k$-sparse parities' using $d^{o(k)}$ queries of tolerance $1/\text{poly}(d, 2^k)$. This contradicts the $d^{\Omega(k)}$ lower bound on the number of queries required to solve $k$-SPLN of tolerance $1/\text{poly}(d, 2^k)$ we get from Theorem~\ref{thm:sq-szorenyi-proof}. 
\end{proof} 

\noindent\textbf{Remark}: Note that this implies there is no $d^{o(1/\epsilon)}$-time gradient descent algorithm that can agnostically learn $\relu(\vw \cdot \vx)$, under the reasonable assumption that for every $i$ the gradients of $\E_{(\vx, y) \sim \D}\left[\left(\relu_{\vw}(\nu(\vx)_{-i}) - \frac{y + 1}{2}\right)^2\right]$ can be computed by $O(d)$ queries whose norms are polynomially bounded. 

\section{Approximation Algorithm} \label{sec:approx}
In this section we give a learning algorithm that runs in polynomial time in all input parameters and outputs a ReLU that has error $O(\opt^{2/3}) + \epsilon$ where $\opt$ is the error of the best-fitting ReLU. The main reduction is a hard thresholding of the labels to create a training set with Boolean labels. We then apply a recent result giving a polynomial-time approximation algorithm for agnostically learning halfspaces over the Gaussian distribution due to Awasthi et. al.~\cite{awasthi2014power}. We present our algorithm and give a proof of its correctness.
\begin{algorithm}[H] 
	\caption{}\label{algo:opt_2/3}
	\hspace*{\algorithmicindent} \textbf{Input} Training set $\mathcal{S}$ of $m$ samples $(\vx^{i}, y^{i})_{i=1}^{m}$, the agnostic halfspace learning algorithm\\
	\hspace*{\algorithmicindent} \quad \quad \quad $\mathcal{A}$ from \cite{awasthi2014power} and a parameter $\alpha$ \\
	\hspace*{\algorithmicindent} \textbf{Output} Weight vector $\widehat{\vw}$
	\begin{algorithmic}[1]
		\State  Construct $S' := \{ (\vx, \sgn(y - \alpha)) \mid (\vx, y) \in S \}.$
		\State Run $\mathcal{A}$ to recover $\widehat{\vw}$ close in $\err_{0/1}$. 
		\State Return $\widehat{\vw}$
	\end{algorithmic}
\end{algorithm}

\begin{theorem}\label{thm:main-thm-upper-bound}
	There is an algorithm (Algorithm~\ref{algo:opt_2/3}) that given $O(\mathsf{poly}(d,1/\epsilon))$ samples $(\vx, y)$ such that $\vx$ is drawn from $N(0, I_{d})$ and $y \in [0, 1]$ recovers a unit vector $\vw$ such that $\err(\relu_{\vw})  \leq O(\opt^{2/3}) + \epsilon$ where $\opt :=  \min_{\|\vw\| = 1} \err(\relu_{\vw}).$
\end{theorem}
\begin{proof} 
Let $\vw^* = \argmin_{\|\vw\| = 1} \err(\relu_\vw)$ and so,  $\err(\relu_{\vw^*}) = \opt$. Define the $S_{good}$ to be the set of points that are $\alpha$-close to the optimal $\relu$, i.e. $S_{good} = \{x: |y - \relu_{\vw^*}(\vx)| \le \alpha\}$. By Markov's inequality,
\[
\Pr[\vx \not \in S_{good}] = \Pr[|y - \relu_{\vw^*}(\vx)| \ge \alpha] \le \frac{\opt}{\alpha^2}.
\]
This implies that all but an $\frac{\opt}{\alpha^2}$ fraction of the points are $\alpha$-close to their corresponding $y$'s. In the first step of Algorithm~\ref{algo:opt_2/3}, the labels become Boolean. Define the $0/1$ error of the vector $\vw$ as follows, $\err_{0/1}(\vw) = \E[\sgn(y - \alpha) \neq \sgn(\vw \cdot \vx)]$. Let $\vw^{\dagger}$ be the argmin of $\err_{0/1}(\vw)$ over all vectors $\vw$ with $\|\vw\|_2 \leq 1$.   Since for all elements in $S_{good} \backslash \{\vv: \vw^* \cdot \vv \in (0, 2 \alpha)\}$, $\sgn(y - \alpha) =  \sgn(\vw^* \cdot \vx)$, 
\begin{align*}
\err_{0/1}(w^*) &\le \Pr[\vx \not\in S_{good} \backslash \{\vv: \vw^* \cdot \vv \in (0, 2 \alpha)\}]\\
& \le \Pr[\vx \not \in S_{good}] + \Pr[\vx \in \{\vv: \vw^* \cdot \vv \in (0, 2 \alpha)\}] \\
& \le \frac{\opt}{\alpha^2} + \frac{1}{\sqrt{2\pi}}\int_0^{2 \alpha} e^{-g^2/2} dg~\le ~\frac{\opt}{\alpha^2} + 2\alpha.
\end{align*}
We now apply Theorem 8 from \cite{awasthi2014power} which gives an algorithm with polynomial running time in $d$ and $1/\epsilon$ that outputs a $\vw$ such that $\| \vw \| = 1$ and $\| \vw - \vw^{\dagger} \| \leq O(\left(\frac{\opt}{\alpha^2} + 2\alpha\right)) + \epsilon$.  For unit vectors $\textbf{a}, \textbf{b}$, $\theta(\textbf{a}, \textbf{b}) < C \Pr[ \sgn(\textbf{a} \cdot \vx) \neq \sgn(\textbf{b} \cdot \vx)]$ for some absolute constant $C$ where $\theta(\textbf{a}, \textbf{b})$ is the angle between the vectors (see Lemma 2 in \cite{awasthi2014power}). The triangle inequality and the fact that $  \| \textbf{a} - \textbf{b} \| \leq \theta(\textbf{a}, \textbf{b})$ implies that if $\err_{0/1}(\textbf{a}), \err_{0/1}(\textbf{b}) < \eta$ then $\|\textbf{a} - \textbf{b}\| \leq C \Pr[ \sgn(\textbf{a} \cdot \vx) \neq \sgn(\textbf{b} \cdot \vx)] \leq O(\eta)$. 
Applying this to $\vw^{\dagger}$ and $\vw^*$ yields $\|{\vw}^{\dagger} - \vw^*\| < O( \frac{\opt}{\alpha^2} + 2\alpha)$. Since the ReLU function is 1-Lipschitz, we have
\begin{align*}
\err(\relu_\vw) &= \E[(y - \relu(\vw \cdot \vx))^2]\\
&\le 2 \E[(y - \relu(\vw^* \cdot \vx))^2] + 2 \E[(\relu(\vw^* \cdot \vx) - \relu(\vw \cdot \vx))^2]\\
&\le 2 \opt + 2 \E[((\vw^* - \vw) \cdot \vx)^2]\\
& = 2 \opt + 2 \|\vw^* - \vw\|^2 ~\le~ O\left( \opt + \left(\frac{\opt}{\alpha^2} + 2\alpha\right)^2 + \epsilon\right)
\end{align*}
Setting $\alpha = \opt^{1/3}$ and rescaling $\epsilon$ we have $\err(\relu_\vw) \le O(\opt^{2/3}) + \epsilon.$
\end{proof}

\section{Conclusions and Open Problems}
We have shown hardness for solving the empirical risk minimization
problem for just one ReLU with respect to
Gaussian distributions and given the first nontrivial approximation
algorithm.  Can we achieve approximation $O(\opt) + \epsilon$?  Note our results holds only for the case of unbiased ReLUs, as the
constant function $1/2$ may achieve smaller square-loss than any
unbiased ReLU.  Interestingly, all positive results that we are aware
of for learning ReLUs (or one-layer ReLU networks) with respect to
Gaussians also assume the ReLU activations are unbiased (e.g.,
\cite{brutzkus2017globally, soltanolkotabi2017learning, conf/icml/GoelKM18, journals/corr/abs-1810-06793,
  journals/corr/abs-1711-00501, journals/corr/abs-1806-07808}).  How
difficult is the biased case?

\bibliographystyle{alpha}
\bibliography{references}
\appendix
\section{Useful Properties}
\begin{fact}[\cite{buet2015multinomial}]\label{fact:Hermite_multinomial}
  For $\beta_1, \ldots, \beta_k$ such that $\sum_{i=1}^k \beta_i^2 = 1$, we have
  \[
  H_n\left(\sum_{i = 1}^{k} \beta_i x_i\right) = \sum_{n_1 + \ldots + n_k = n}\frac{n!}{n_1! \cdots n_k!}\prod_{j=1}^k \beta_j^{n_j}H_{n_j}(x_j).
  \]
\end{fact}

\begin{lemma}\label{lem:reluexp}
For any $\vw$, we have
\[
\E_{\vz \sim \mathcal{N}(0, \rmI_{d})}\left[\relu(\vw \cdot \vz)^2\right] = \frac{\|\vw\|^2}{2} \text{ and } \E_{\vz \sim \mathcal{N}(0, \rmI_{d})}\left[\relu(\vw \cdot \vz)\right] = \frac{\|\vw\|}{\sqrt{2\pi}}.
\]
\end{lemma}
\begin{proof}
Observe that for $\vz \sim \mathcal{N}(0, \rmI_{d})$, $\vw \cdot \vz \sim \mathcal{N}(0, ||\vw||^2)$, thus we have
\begin{align*}
\E_{\vz \sim \mathcal{N}(0, \rmI_{d})}\left[\relu(\vw \cdot \vz)^2\right]&= \E_{g \sim \mathcal{N}(0, ||\vw||^2)}\left[\relu(g)^2\right]\\
&= \frac{1}{\sqrt{2 \pi}||\vw||} \int_{0}^\infty g^2 e^{- \frac{g^2}{2||\vw||^2}} dg\\
&= \frac{||\vw||^2}{2}
\end{align*}
The last follows from observing that the integral is $1/2$ of the variance of a $N(0, ||\vw||^2)$ variable. Similarly, we have
\begin{align*}
\E_{\vz \sim \mathcal{N}(0, \rmI_{d})}\left[\relu(\vw \cdot \vz)\right]&= \E_{g \sim \mathcal{N}(0, ||\vw||^2)}\left[\relu(g)\right]\\
&= \frac{1}{\sqrt{2 \pi}||\vw||} \int_{0}^\infty g e^{- \frac{g^2}{2||\vw||^2}} dg\\
&= \frac{||\vw||}{\sqrt{2\pi}}.
\end{align*}
Here the last equality follows from standard computation of mean of the absolute value of a Gaussian random variable.
\end{proof}

\section{Omitted Proofs}

\begin{proof}[Proof of Claim~\ref{lem:hrelu}]
  For $i \ge 2$,
  \begin{align}
  \hat{\relu}_i &= \frac{1}{\sqrt{2 \pi}}\int_{-\infty}^{\infty} \relu(x) H_i(x) e^{-\frac{x^2}{2}} dx\\
  &= \frac{1}{\sqrt{2 \pi i!}}\int_{0}^{\infty} xH_i(x) e^{-\frac{x^2}{2}} dx \\
  &= \frac{1}{\sqrt{2 \pi i!}}\int_{0}^{\infty} (H_{i + 1}(x) + iH_{i-1}(x)) e^{-\frac{x^2}{2}} dx \\
  &= \frac{1}{\sqrt{2 \pi i!}} (H_{i}(0) + iH_{i-2}(0))
  \end{align}
  Here we used the additional property on the recurrence of $H$, that is, $H_{n+1}(x) = xH_{n}(x) - nH_{n-1}$.
\end{proof}

\begin{proof}[Proof of Claim~\ref{claim:h_expansion_relu}] 
    Since $\sum_{i \in S} \left(\frac{1}{\sqrt{k}}\right)^2 = 1$ using Fact \ref{fact:Hermite_multinomial}, we have
    \begin{align}
    \relu\left(\frac{\sum_{i \in S} z_i}{\sqrt{k}}\right) &= \sum_{n=0}^{\infty} \frac{\hat{\relu}_n}{\sqrt{n!}} \cdot H_n\left(\frac{\sum_{i \in S} z_i}{\sqrt{k}}\right) \\
    &= \sum_{n=0}^{\infty} \frac{\hat{\relu}_n}{\sqrt{n!}} \cdot\left[ \sum_{n_1 + \ldots + n_k = n}\frac{n!}{n_1! \cdots n_k!}\prod_{j=1}^k \frac{1}{k^{n_j/2}} H_{n_j}(z_j) \right] \\
    &= \sum_{n=0}^{\infty} \frac{\hat{\relu}_n}{\sqrt{n!}} \cdot\left[ \frac{1}{k^{n/2}} \cdot \sum_{n_1 + \ldots + n_k = n}\frac{n!}{n_1! \cdots n_k!}\prod_{j=1}^k H_{n_j}(z_j) \right] \\
    &=\sum_{n=0}^{\infty} \frac{\hat{\relu}_n}{\sqrt{n!}} \cdot \left[ \frac{1}{k^{n/2}}\sum_{n_1 + \ldots + n_k = n}\frac{n!}{n_1! \cdots n_k!}\prod_{j=1}^k \sqrt{n_j!}\bar{H}_{n_j}(z_j) \right]\\
    &= \sum_{n=0}^{\infty} \frac{\hat{\relu}_n}{k^{n/2}}\cdot\sum_{n_1 + \ldots + n_k = n}\left(\frac{n!}{n_1! \cdots n_k!}\right)^{1/2}\prod_{j=1}^k\bar{H}_{n_j}(z_j)
    \end{align}
  \end{proof}

\begin{proof}[Proof of Claim~\ref{claim:inactive_index}]
Since $j \in S$, removing index $j$, $\E_{z_j}[\chi_S(\vz)|\vz_{-j}] = 0$. Thus, for the input, the label is a Bernoulli random variable with probability $1/2$. Thus we have,
\begin{align}
&\err_{D_j}(\relu_\vw)\\
&= \E_{\vz \sim \mathcal{N}(0, \rmI_d)}[(\relu(\vw \cdot \vz_{-j}) - y^\prime)^2]\\
& = \frac{1}{2}\left(\E_{\vz_{-j} \sim \mathcal{N}(0, \rmI_{d-1})}\left[\left(\relu(\vw \cdot \vz_{-j}) - 1\right)^2\right] + \E_{\vz_{-j} \sim \mathcal{N}(0, \rmI_{d-1})}\left[\left(\relu(\vw \cdot \vz_{-j})\right)^2\right]\right)\\
&= \E_{\vz_{-j} \sim \mathcal{N}(0, \rmI_{d-1})}\left[\relu(\vw \cdot \vz_{-j})^2\right] - \E_{\vz_{-j} \sim \mathcal{N}(0, \rmI_{d-1})}\left[\relu(\vw \cdot \vz_{-j})\right] + \frac{1}{2}\\
&= \frac{||\vw||^2}{2} - \frac{||\vw||}{\sqrt{2\pi}} + \frac{1}{2}
\end{align}
Here the third equality follows since $j \in S$ and not in $\vz_{-j}$ therefore, the label is random for the ReLU. The last equality follows from Lemma~\ref{lem:reluexp}. Note that, for any ReLU, the minimum error is achieved when $||\vw|| = \frac{1}{\sqrt{2\pi}}$. Thus when $j \not \in S$ the best ReLU achieves error at least $\frac{1}{2} - \frac{1}{4\pi}$.
  \end{proof}

\begin{proof}[Proof of Claim~\ref{claim:active_index}]
Since $j$ is not a relevant variable $S \subseteq [d] \setminus \{ j \}$, from Theorem \ref{thm:relu_correlated}, we know that there exists $\relu_{\vw_S}$ with $||\vw_S|| = 1/\sqrt{2\pi}$ dependent only on variables in $S$ correlated with $\chi^\gamma_S$,
  \begin{align}
  &\err_{D_j}(\relu_{\vw_S})\\
    &= \E_{\vz \sim \mathcal{N}(0, \rmI_{d})}[(\relu(\vw_S \cdot \vz) - y^\prime)^2]\\
  &= (1- \eta) \E_{\vz \sim \mathcal{N}(0, \rmI_d)}\left[\left(\relu(\vw_S \cdot \vz) - \frac{\prod_{i \in S}\sgn(z_i) + 1}{2}\right)^2\right] \\
    &\qquad + \eta \E_{\vz \sim \mathcal{N}(0, \rmI_d)}\left[\left(\relu(\vw_S \cdot \vz) - \frac{1 - \prod_{i \in S}\sgn(z_i)}{2}\right)^2\right]\\
  &= \E_{\vz \sim \mathcal{N}(0, \rmI_{d})}\left[\relu(\vw_S \cdot \vz_S)^2\right] - (1 - 2\eta)\E_{\vz \sim \mathcal{N}(0, \rmI_{d})}\left[\relu(\vw_S \cdot \vz)\prod_{i \in S}\sgn(z_i)\right]\\
  &\quad \quad - \E_{\vz \sim \mathcal{N}(0, \rmI)}\left[\relu(\vw_S \cdot \vz)\right]  +  \frac{1}{2}\E_{\vz \sim \mathcal{N}(0, \rmI_{d})}\left[\prod_{i \in S}\sgn(z_i)^2\right]\\
  &= \frac{||\vw_S||^2}{2} -\frac{||\vw_S||}{\sqrt{2 \pi}} + \frac{1}{2} - (1 - 2\eta)\E_{\vz \sim \mathcal{N}(0, \rmI_{d})}\left[\relu(\vw_S \cdot \vz)\prod_{i \in S}\sgn(z_i)\right]\\
  &\ge \frac{||\vw^*||^2}{2} -\frac{||\vw^*||}{\sqrt{2 \pi}} + \frac{1}{2} - \frac{2^{-O(k)}}{1 - 2 \eta} = \frac{1}{2} - \frac{1}{4\pi} - \frac{2^{-O(k)}}{1 - 2 \eta}
  \end{align}
\end{proof}

\end{document}